Hello!
I am trying to solve the following for homework. I am really struggling with functions.
Sketch the graph of the function: f(x)=(2/3)^x
Thanks so much for the help!
Jim

Answer by rfer(16322)

f(2)=(2/3)^2
f(2)=(2/3)(2/3)
f(2)=4/9
--------------------
f(3)=(2/3)^3
f(3)=(2/3)(2/3)(2/3)
f(3)=8/27